OpenStudy (anonymous):

divide or multiply then simplify. 1) square root of 8y^5 * square root of 40y^2 2) square root of 48a^5b divided by square root of 12ab

OpenStudy (anonymous):

square root of 8y^5 *square root of 40y^2 = square root of (8y^5 * 40y^2) = square root of ( 8^2 *5 (y^3)^2 *y) =8y^3 square root of (5y)

OpenStudy (anonymous):

any confusion

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

which step

OpenStudy (anonymous):

everything?

OpenStudy (anonymous):

ok see second part

OpenStudy (anonymous):

(squareroot(48*a^5*b))/(sqrt(12ab)) now sqrt(48)=4*sqrt(3) sqrt(12)=2*sqrt(3) is it right??fine till here